Mohan can solve 90% of the problems of his mathematics book while sohan can solve 70% of them. If a problem is selected at ramdom, what is the probability that none of them will solve the problem?

Grade:12

1 Answers

Ashwin Muralidharan IIT Madras
12 years ago

Hi Shaleen,

 

The Probability that none of them solve is:

 

P(Mohan doesn't solve & Sohan doesn't solve) = 1/10 * 3/10 = 3/100 = 3% chance that none of them solve it.

 

Hope that helps.

Best Regards,

Ashwin (IIT MadraS).